# Multi-Objective Agentic Rewrites for Unstructured Data Processing

### v1

@@ 5.2.1. Accuracy Improvement
Table 4 summarizes MOAR’s accuracy improvements across all baselines. MOAR achieves the highest accuracy on every workload, with gains over the next most accurate baseline ranging from +2.2% (Medec, Sustainability) to +415.5% (BlackVault). The smallest accuracy improvement is on Medec (+2.2% over the Simple Agent), where the Simple Agent’s top-accuracy pipeline simply replaced the default model with gpt-5. One possible explanation is that documents are very short (147 words on average), and we use the optimized prompt from the original paper (abacha2024medec)—leaving minimal room for further optimization.
@@ 5.3. Insights from MOAR’s Pipelines
  • Optimal models are workload-dependent. All 29 pipelines switch from the default gpt-4o-mini, even though it is OpenAI’s “most cost-efficient small model” (openai_gpt4o_mini). Each workload’s most accurate pipeline uses a different model. The model usage patterns reveal specializations (Table 6): gpt-5-nano is the most prevalent (41%), used for extraction and classification on both long and short documents, while smaller models like gemini-2.5-flash-lite and gpt-4o-mini are used for summarization on long documents.
@@ 6  Related Work
Semantic Data Processing Systems. Systems that expose AI-powered data processing capabilities have proliferated over the past two years. ThalamusDB (jo2024thalamusdb) was among the earliest to support natural language filters and joins in SQL. LOTUS (patel2025semantic) coined the term “semantic operators” for this type of LLM-powered data processing; other systems that support at least one semantic operator include Palimpzest (liu2024declarative), Aryn (anderson2024design), DocETL (shankar2025docetl), FlockMTL (flockdb), ELEET (urban2024eleet), Unify (unify), and DocDB (docdb). Other systems target specific applications like templatized documents (lin2025querying), extraction tasks (sun2025quest; arora2023language), and social science queries (hu2024leap). Industrial systems from Databricks (databricks-llm), DuckDB (duckdb-llm), Snowflake (snowflake-llm; aggarwal2025cortex), Google BigQuery (HeSethi2025BigQueryAIFunctions), and Google AlloyDB (alloydb-llm) also now support LLM-powered functions. In our evaluation (Sec. 5), we test active open-source systems that support at least a semantic map operator—LOTUS, Palimpzest, and DocETL-V1.
@@
abacus (russo2025abacus), flock-mtl (flockdb), unify (unify), docdb (docdb), moar
@@
docetl-v1 (shankar2025docetl), caesura (urban2024demonstrating), unify (unify), aop (wang2025aop), moar
@@
abacus (russo2025abacus), lotus (patel2025semantic), thalamusdb (jo2024thalamusdb), cortex-aisql (aggarwal2025cortex), eleet (urban2024eleet), unify (unify), docdb (docdb), moar
@@
docetl-v1 (shankar2025docetl), cortex-aisql (aggarwal2025cortex), moar
@@
    Plan space. Classical relational database optimizers search over rewrites such as filter and join reordering (chaudhuri1998overview; hellerstein1993predicate; selinger1979access). These are rule-based, algebraic transformations. Semantic operators admit a richer space of rewrites—including transformations over the meaning of the task and data—which we organize along two axes (Fig. 5): whether rewrites are data-dependent (requiring samples to instantiate) or data-independent, and whether rewrites are agentic (synthesized by an LLM, as in DocETL) or rule-based (not requiring an LLM agent to instantiate). Examples of rule-based, data-independent rewrites include physical implementations like model substitution and ensembling in ABACUS (russo2025abacus). Examples of rule-based, data-dependent rewrites—those that require samples to learn configurations but not an LLM agent to instantiate—include model cascades (patel2025semantic; viola2001rapid; kang2017noscope; zeighami2025cut), context reduction (russo2025abacus), and fine-tuning small LLMs (urban2024eleet). An example of a rewrite that requires an LLM agent to instantiate, but not sample data, is operator fusion. Examples of agentic, data-dependent rewrites include most MOAR rewrite directives. These quadrants vary in instantiation cost—from cheap (rule-based, data-independent) to expensive (agentic, data-dependent)—but also in expressiveness: e.g., agentic rewrites can synthesize novel transformations. MOAR is the only query optimizer with rewrites that span all four quadrants.
@@
    Cost model. A cost model estimates the quality of candidate plans, guiding the search algorithm toward good plans. Selinger’s cost model, for example, estimates I/O and CPU costs using statistics like cardinality and index availability (selinger1979access). For semantic operators, the new challenge is estimating accuracy in addition to monetary cost or latency: there is no closed-form expression for how well an LLM will perform a task (semanticic; spade). Approaches to accuracy estimation vary in cost and precision. The simplest approach naively executes candidate plans on samples (as in LOTUS, DocETL-V1, and MOAR). ABACUS uses multi-armed bandits to adaptively sample plans and estimate their accuracy, tightening the accuracy bounds for promising (i.e., Pareto frontier) candidate plans; this cost estimation technique could be integrated into MOAR’s search algorithm. Some rewrites guarantee accuracy by construction: model cascades and approximate query processing techniques bound accuracy relative to the unrewritten plan (patel2025semantic; jo2024thalamusdb; jo2024smart; zeighami2025cut). However, all approaches require some notion of ground truth—either a user-specified “oracle” LLM-powered implementation (patel2025semantic; russo2025abacus; jo2024smart), a user-defined accuracy function (russo2025abacus), or LLM-as-judge pairwise comparisons for ranking candidate plans (shankar2025docetl).
@@
    Search algorithm. Given a plan space and cost model, a search algorithm finds good plans. Classical optimizers use dynamic programming, composing optimal subplans into optimal plans (graefe1993volcano; Graefe1995TheCF). For semantic operators, two new challenges arise: plans may lack optimal substructure (as explained in Sec. 1), and with accuracy as an objective, optimizers should return a Pareto frontier of plans, so users can choose their preferred cost-accuracy tradeoff. Only ABACUS, DocETL-V1, and MOAR search over a space of rewrites: ABACUS adapts Cascades (Graefe1995TheCF) to return a Pareto frontier; DocETL-V1 introduces a top-down search algorithm designed for LLM-as-judge evaluation; MOAR uses UCT-based search (kocsis2006bandit) over complete pipelines, avoiding optimal substructure assumptions. ThalamusDB and LOTUS search for optimal parameters within specific rewrites (e.g., model cascade thresholds) but do not search across different rewrites. Other solutions provide interactive interfaces to help users rewrite pipelines themselves (liu2025palimpchat; shankar2025steering).
@@ 7  Conclusion
We introduce MOAR (Multi-Objective Agentic Rewrites), a novel optimizer for LLM-powered data processing pipelines that jointly optimizes for both accuracy and cost. Building on DocETL’s foundation of agentic rewrite directives, MOAR introduces (i) an expanded library of over 30 rewrite directives, (ii) a multi-armed bandit-based search algorithm that efficiently discovers sequences of rewrites that lead to good plans, and (iii) comprehensive empirical validation across six real-world workloads, demonstrating substantial improvements over state-of-the-art systems. MOAR achieves the highest accuracy on all workloads. Compared to ABACUS (russo2025abacus), the next-best optimizer, MOAR achieves 27% higher accuracy on average while matching its best accuracy at only 55% of its cost.
@@ Appendix B Detailed Descriptions of New Rewrite Directives
Throughout this section, we adopt the notation from shankar2025docetl: given operators $A$ and $B$, we denote their composition as $A\to B$, where $(A\to B)(D)=B(A(D))$. For independent execution, we use $A\parallel B$. We may drop arguments (e.g., $\text{Map}_{x}(D)$ becomes $\text{Map}_{x}$) and omit subscripts except when the same operator appears multiple times. We color new or modified operators introduced by a rewrite in green. The arrow $\Rightarrow$ denotes a rewrite of the operator (or operator sequence) on the left into the form on the right.
@@ B.1.5. Reordering
MOAR applies this rewrite only when the LLM agent verifies that reordering preserves pipeline semantics, e.g., the rewritten filter does not depend on attributes produced by the operator it is moved before. In practice, user-authored pipelines are short (2–3 operators) (shankar2025steering), so reordering becomes more valuable as MOAR’s search generates longer, more complex pipelines through sequences of rewrites.
@@ B.3.1. Chunk Sampling
The Sample operator can be instantiated with random sampling, keyword search (based on BM25 retrieval (robertson2009probabilistic)), or embedding-based similarity, selecting the top-$k$ document chunks relevant to a query (lewis2020retrieval). When instantiating the directive, the LLM agent synthesizes the sampling method, the query (if not random sampling), and the value of $k$.
@@ B.5. LLM-Centric Rewrites
DocETL introduced directives that improve LLM output quality by refining how tasks are specified to the LLM. For example, the gleaning directive uses a validator LLM to check outputs and provide feedback for iterative refinement. Similar strategies have been explored in ABACUS, which implements a “critique-and-refine” physical implementation of map operations (russo2025abacus). Inspired by prompting strategies (liuprompting2023) and optimization techniques (ramnath2025systematic), MOAR adds directives that improve prompt quality and provide examples to guide LLM behavior.
@@ B.5.1. Model Substitution
where $x=(t,s,m)$ and $x^{\prime}=(t,s,m^{\prime})$ with $m^{\prime}\neq m$. When instantiating this directive, the LLM agent receives context about model performance: for each model in the available pool $M$, the agent sees the cost and accuracy achieved by the original pipeline when executed with that model on a sample of data. The agent also has access to each model’s performance on MRCR (a long-context benchmark that evaluates an LLM’s ability to retrieve and distinguish between multiple similar requests hidden in long contexts (openai_mrcr; vodrahalli2024michelangelo)), as well as context window size and pricing details.
@@ B.5.3. Find Few-shot Examples
A popular prompt engineering technique is to include few-shot examples in the prompt to demonstrate desired behavior (brown2020language). This directive synthesizes such examples to improve operator accuracy:
@@
where $x^{\prime}=(t^{\prime},s,m)$ and $t^{\prime}$ is a modified prompt template that incorporates few-shot examples. The agent examines sample documents, generates input-output pairs that demonstrate the desired behavior, and constructs the modified prompt with these examples embedded. Note that tools like DSPy (khattab2024dspy) could automate few-shot example generation when this directive is chosen. However, DSPy’s iterative optimization process (which evaluates multiple example sets to find the best one) takes significantly longer than our agent’s “single-pass” instantiation. Supporting such iterative optimization would require modifying our search algorithm to account for varying directive instantiation costs—we leave this to future work.
@@ B.5.4. Arbitrary Rewrite
To implement arbitrary rewrites, we pass the entire pipeline as YAML code to the agent and ask it to propose edits using a find-and-replace approach inspired by how coding agents work (aider_edit_formats). Specifically, the agent returns a list of search-and-replace blocks, where each block specifies: (i) a unique string to search for in the original pipeline YAML, and (ii) the replacement text. After the agent produces an arbitrary rewrite, we verify that the resulting pipeline can be parsed and executed by DocETL. If parsing or validation fails, we provide the error message to the agent and retry up to 3 times before discarding the rewrite.

### v3

@@ 5.2.1. Accuracy Improvement
Table 4 summarizes MOAR’s accuracy improvements across all baselines. MOAR achieves the highest accuracy on every workload, with gains over the next most accurate baseline ranging from +2.2% (Medec, Sustainability) to +415.5% (BlackVault). The smallest accuracy improvement is on Medec (+2.2% over the Simple Agent), where the Simple Agent’s top-accuracy pipeline simply replaced the default model with gpt-5. One possible explanation is that documents are very short (147 words on average), and we use the optimized prompt from the original paper (Abacha et al., 2024)—leaving minimal room for further optimization.
@@ 5.3. Insights from MOAR’s Pipelines
  • Optimal models are workload-dependent. All 29 pipelines switch from the default gpt-4o-mini, even though it is OpenAI’s “most cost-efficient small model” (OpenAI, 2024). Each workload’s most accurate pipeline uses a different model. The model usage patterns reveal specializations (Table 6): gpt-5-nano is the most prevalent (41%), used for extraction and classification on both long and short documents, while smaller models like gemini-2.5-flash-lite and gpt-4o-mini are used for summarization on long documents.
@@ 6  Related Work
Semantic Data Processing Systems. Systems that expose AI-powered data processing capabilities have proliferated over the past two years. ThalamusDB (Jo and Trummer, 2024) was among the earliest to support natural language filters and joins in SQL. LOTUS (Patel et al., 2025) coined the term “semantic operators” for this type of LLM-powered data processing; other systems that support at least one semantic operator include Palimpzest (Liu et al., 2025a), Aryn (Anderson et al., 2025), DocETL (Shankar et al., 2025a), FlockMTL (Dorbani et al., 2025), ELEET (Urban and Binnig, 2024b), Unify (Wang et al., 2025), and DocDB (Li et al., 2025). Other systems target specific applications like templatized documents (Lin et al., 2025), extraction tasks (Sun et al., 2025; Arora et al., 2023), and social science queries (Hu et al., 2024). Industrial systems from Databricks (Wendell et al., 2023), DuckDB (Döhmen, 2024), Snowflake (Snowflake, 2025; Aggarwal et al., 2025), Google BigQuery (He and Sethi, 2025), and Google AlloyDB (Google Cloud, 2025) also now support LLM-powered functions. In our evaluation (Sec. 5), we test active open-source systems that support at least a semantic map operator—LOTUS, Palimpzest, and DocETL-V1.
@@
abacus (Russo et al., 2025), flock-mtl (Dorbani et al., 2025), unify (Wang et al., 2025), docdb (Li et al., 2025), moar
@@
docetl-v1 (Shankar et al., 2025a), caesura (Urban and Binnig, 2024a), unify (Wang et al., 2025), aop (Wang and Li, 2025), moar
@@
abacus (Russo et al., 2025), lotus (Patel et al., 2025), thalamusdb (Jo and Trummer, 2024), cortex-aisql (Aggarwal et al., 2025), eleet (Urban and Binnig, 2024b), unify (Wang et al., 2025), docdb (Li et al., 2025), moar
@@
docetl-v1 (Shankar et al., 2025a), cortex-aisql (Aggarwal et al., 2025), moar
@@
    Plan space. Classical relational database optimizers search over rewrites such as filter and join reordering (Chaudhuri, 1998; Hellerstein and Stonebraker, 1993; Selinger et al., 1979). These are rule-based, algebraic transformations. Semantic operators admit a richer space of rewrites—including transformations over the meaning of the task and data—which we organize along two axes (Fig. 5): whether rewrites are data-dependent (requiring samples to instantiate) or data-independent, and whether rewrites are agentic (synthesized by an LLM, as in DocETL) or rule-based (not requiring an LLM agent to instantiate). Examples of rule-based, data-independent rewrites include physical implementations like model substitution and ensembling in ABACUS (Russo et al., 2025). Examples of rule-based, data-dependent rewrites—those that require samples to learn configurations but not an LLM agent to instantiate—include model cascades (Patel et al., 2025; Viola and Jones, 2001; Kang et al., 2017; Zeighami et al., 2025), context reduction (Russo et al., 2025), and fine-tuning small LLMs (Urban and Binnig, 2024b). An example of a rewrite that requires an LLM agent to instantiate, but not sample data, is operator fusion. Examples of agentic, data-dependent rewrites include most MOAR rewrite directives. These quadrants vary in instantiation cost—from cheap (rule-based, data-independent) to expensive (agentic, data-dependent)—but also in expressiveness: e.g., agentic rewrites can synthesize novel transformations. MOAR is the only query optimizer with rewrites that span all four quadrants.
@@
    Cost model. A cost model estimates the quality of candidate plans, guiding the search algorithm toward good plans. Selinger’s cost model, for example, estimates I/O and CPU costs using statistics like cardinality and index availability (Selinger et al., 1979). For semantic operators, the new challenge is estimating accuracy in addition to monetary cost or latency: there is no closed-form expression for how well an LLM will perform a task (Lee et al., 2025; Shankar et al., 2024). Approaches to accuracy estimation vary in cost and precision. The simplest approach naively executes candidate plans on samples (as in LOTUS, DocETL-V1, and MOAR). ABACUS uses multi-armed bandits to adaptively sample plans and estimate their accuracy, tightening the accuracy bounds for promising (i.e., Pareto frontier) candidate plans; this cost estimation technique could be integrated into MOAR’s search algorithm. Some rewrites guarantee accuracy by construction: model cascades and approximate query processing techniques bound accuracy relative to the unrewritten plan (Patel et al., 2025; Jo and Trummer, 2024, 2025; Zeighami et al., 2025). However, all approaches require some notion of ground truth—either a user-specified “oracle” LLM-powered implementation (Patel et al., 2025; Russo et al., 2025; Jo and Trummer, 2025), a user-defined accuracy function (Russo et al., 2025), or LLM-as-judge pairwise comparisons for ranking candidate plans (Shankar et al., 2025a).
@@
    Search algorithm. Given a plan space and cost model, a search algorithm finds good plans. Classical optimizers use dynamic programming, composing optimal subplans into optimal plans (Graefe and McKenna, 1993; Graefe, 1995). For semantic operators, two new challenges arise: plans may lack optimal substructure (as explained in Sec. 1), and with accuracy as an objective, optimizers should return a Pareto frontier of plans, so users can choose their preferred cost-accuracy tradeoff. Only ABACUS, DocETL-V1, and MOAR search over a space of rewrites: ABACUS adapts Cascades (Graefe, 1995) to return a Pareto frontier; DocETL-V1 introduces a top-down search algorithm designed for LLM-as-judge evaluation; MOAR uses UCT-based search (Kocsis and Szepesvári, 2006) over complete pipelines, avoiding optimal substructure assumptions. ThalamusDB and LOTUS search for optimal parameters within specific rewrites (e.g., model cascade thresholds) but do not search across different rewrites. Other solutions provide interactive interfaces to help users rewrite pipelines themselves (Liu et al., 2025b; Shankar et al., 2025b).
@@ 7  Conclusion
We introduce MOAR (Multi-Objective Agentic Rewrites), a novel optimizer for LLM-powered data processing pipelines that jointly optimizes for both accuracy and cost. Building on DocETL’s foundation of agentic rewrite directives, MOAR introduces (i) an expanded library of over 30 rewrite directives, (ii) a multi-armed bandit-based search algorithm that efficiently discovers sequences of rewrites that lead to good plans, and (iii) comprehensive empirical validation across six real-world workloads, demonstrating substantial improvements over state-of-the-art systems. MOAR achieves the highest accuracy on all workloads. Compared to ABACUS (Russo et al., 2025), the next-best optimizer, MOAR achieves 27% higher accuracy on average while matching its best accuracy at only 55% of its cost.
@@ Appendix B Detailed Descriptions of New Rewrite Directives
Throughout this section, we adopt the notation from Shankar et al. (2025a): given operators $A$ and $B$, we denote their composition as $A\to B$, where $(A\to B)(D)=B(A(D))$. For independent execution, we use $A\parallel B$. We may drop arguments (e.g., $\text{Map}_{x}(D)$ becomes $\text{Map}_{x}$) and omit subscripts except when the same operator appears multiple times. We color new or modified operators introduced by a rewrite in green. The arrow $\Rightarrow$ denotes a rewrite of the operator (or operator sequence) on the left into the form on the right.
@@ B.1.5. Reordering
MOAR applies this rewrite only when the LLM agent verifies that reordering preserves pipeline semantics, e.g., the rewritten filter does not depend on attributes produced by the operator it is moved before. In practice, user-authored pipelines are short (2–3 operators) (Shankar et al., 2025b), so reordering becomes more valuable as MOAR’s search generates longer, more complex pipelines through sequences of rewrites.
@@ B.3.1. Chunk Sampling
The Sample operator can be instantiated with random sampling, keyword search (based on BM25 retrieval (Robertson et al., 2009)), or embedding-based similarity, selecting the top-$k$ document chunks relevant to a query (Lewis et al., 2020). When instantiating the directive, the LLM agent synthesizes the sampling method, the query (if not random sampling), and the value of $k$.
@@ B.5. LLM-Centric Rewrites
DocETL introduced directives that improve LLM output quality by refining how tasks are specified to the LLM. For example, the gleaning directive uses a validator LLM to check outputs and provide feedback for iterative refinement. Similar strategies have been explored in ABACUS, which implements a “critique-and-refine” physical implementation of map operations (Russo et al., 2025). Inspired by prompting strategies (Liu et al., 2023) and optimization techniques (Ramnath et al., 2025), MOAR adds directives that improve prompt quality and provide examples to guide LLM behavior.
@@ B.5.1. Model Substitution
where $x=(t,s,m)$ and $x^{\prime}=(t,s,m^{\prime})$ with $m^{\prime}\neq m$. When instantiating this directive, the LLM agent receives context about model performance: for each model in the available pool $M$, the agent sees the cost and accuracy achieved by the original pipeline when executed with that model on a sample of data. The agent also has access to each model’s performance on MRCR (a long-context benchmark that evaluates an LLM’s ability to retrieve and distinguish between multiple similar requests hidden in long contexts (OpenAI, ; Vodrahalli et al., 2024)), as well as context window size and pricing details.
@@ B.5.3. Find Few-shot Examples
A popular prompt engineering technique is to include few-shot examples in the prompt to demonstrate desired behavior (Brown et al., 2020). This directive synthesizes such examples to improve operator accuracy:
@@
where $x^{\prime}=(t^{\prime},s,m)$ and $t^{\prime}$ is a modified prompt template that incorporates few-shot examples. The agent examines sample documents, generates input-output pairs that demonstrate the desired behavior, and constructs the modified prompt with these examples embedded. Note that tools like DSPy (Khattab et al., 2024) could automate few-shot example generation when this directive is chosen. However, DSPy’s iterative optimization process (which evaluates multiple example sets to find the best one) takes significantly longer than our agent’s “single-pass” instantiation. Supporting such iterative optimization would require modifying our search algorithm to account for varying directive instantiation costs—we leave this to future work.
@@ B.5.4. Arbitrary Rewrite
To implement arbitrary rewrites, we pass the entire pipeline as YAML code to the agent and ask it to propose edits using a find-and-replace approach inspired by how coding agents work (Aider Project, 2025). Specifically, the agent returns a list of search-and-replace blocks, where each block specifies: (i) a unique string to search for in the original pipeline YAML, and (ii) the replacement text. After the agent produces an arbitrary rewrite, we verify that the resulting pipeline can be parsed and executed by DocETL. If parsing or validation fails, we provide the error message to the agent and retry up to 3 times before discarding the rewrite.